OpenStudy (anonymous):

The plane through (1,2-1) that is perpendicular to the line of intersection of the planes 2x+y+z= 2 and x+2y+z = 3

OpenStudy (anonymous):

Find the equation of the plane

OpenStudy (anonymous):

The thing i don'y understand is that when you take cross product of the two normals to get the desired normal which is (-1,-1,3) why do we take it as (1,1,-3)?

ganeshie8 (ganeshie8):

it doesnt matter, they're direction ratios, so any of (-1k, -1k, 3k) will work

ganeshie8 (ganeshie8):

maybe you're changing it to(1, 1, -3) cuz it has fewer negatives

OpenStudy (anonymous):

but the equation would be different then right?

OpenStudy (anonymous):

Its done that way in the manual

ganeshie8 (ganeshie8):

how, can we walk through and see wat we get, i dont remember exactly how to actually solve this

OpenStudy (amistre64):

i remember :)

OpenStudy (anonymous):

we get the two normals from the equations and then take their cross product. Then use that normal the point given to get the equation

ganeshie8 (ganeshie8):

the intersection of given planes is a line, since this line is perpendicular to the required plane, the equation wud be : a(x-x1)+b(y-y1)+c(z-z1) = 0 since we knw one point, we're done. im thinking we can solve as above

OpenStudy (amistre64):

a plane that is perp to a line, uses the lines direction vector as its normal .... yep

OpenStudy (amistre64):

define 2 points to establish a vector between

OpenStudy (amistre64):

we have a system of equations: 2x+y+z= 2 x+2y+z = 3 if we hold one variable constant, this reduces to a 2x2 system. let x=0, solve for y,z let z=0, solve for x,y we then have 2 points to vector

OpenStudy (anonymous):

yep thats the equation. we want (a,b,c) which we get through the cross product of (2,1,1) and (1,2,1)

OpenStudy (amistre64):

or that .. yeah

OpenStudy (anonymous):

but why take negative of the normal that we get

OpenStudy (anonymous):

The cross product comes out as (-1,-1,3) but the normal is (1,1,-3)

OpenStudy (amistre64):

-(z+2x+2y) x y z x y 2 1 1 2 1 1 2 1 1 2 +(x+y+4z) x +y+4z -2x-2y -z ----------- -1,-1,3 you do realize that we could have also gotten 1,1,-3 if we had swapped rows right?

OpenStudy (amistre64):

by convention, a direction vector (a,b,c) is built such that a>=0, and a,b,c relative primes

OpenStudy (anonymous):

please explain a little more

OpenStudy (amistre64):

by convention ... as a rule of thumb that everyone agrees on ... becuase we like to have some sort of uniformity of results ... becuase this is what people have agreed to use ... etc,etc,etc. a direction vector (a,b,c) is built such that a>=0, and a,b,c relative primes

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

All above show u same direction, (i missed putting arrows in other direction)

OpenStudy (amistre64):

using any form of k(1,1,-3) will produce the same plane for any nonzero k

OpenStudy (amistre64):

asking why we do it is the same as asking why everyone drives on the right side of the road in America .... it is simply something we all have agreed to do to make an otherwise chaotix system organized

ganeshie8 (ganeshie8):